# Neural Fields with Thermal Activations for Arbitrary-Scale Super-Resolution

### v2

@@ Appendix 0.E Replication of Other Methods
We encountered challenges attempting to recreate results of some of the competing methods, which explains why some numbers are missing in various tables or differ from the originally reported ones. Details are provided below.
@@
CUF [43]. At the time of writing, there were no public code or checkpoints available for CUF. Therefore, we could not generate numbers for datasets and scaling factors not reported in the original paper. We have denoted those missing values with “—” in the tables. Furthermore, we could not create any qualitative samples and could not include CUF in the multi-scale consistency study in Sec. 4.2.
@@
CLIT [10]. For CLIT, at the time of writing, there is public code available, however, checkpoints are not. We attempted to recreate their models, but due to the cascaded training schedule and the big model size, the training process would require excessive amounts of GPU time: over a month using 8×8$\times$8 × Nvidia GeForce RTX 3090s. The authors did unfortunately not respond to our requests for the trained checkpoints used in their paper.
@@
CiaoSR [8]. We noticed that in the official CiaoSR implementation, border cropping prior to evaluation on the DIV2K dataset deviated slightly from all other methods. We aligned this to enable a meaningful comparison and re-computed all DIV2K numbers. This explains the slight deviation from the original paper.
DIV2K ×$\times 2$
DIV2K ×$\times 3$
DIV2K ×$\times 3$
Urban100 ×$\times 2$
Urban100 ×$\times 3$
Urban100 ×$\times 3$
Manga109 ×$\times 3$
Manga109 ×$\times 3$
LR input LIIF [11] LTE [22] CiaoSR [8] Thera-L (ours) GT
Figure 9: Further qualitative examples for ×$\times 2$ and ×$\times 3$ upsampling factors. Best viewed zoomed in.
DIV2K ×$\times 4$
DIV2K ×$\times 4$
Urban100 ×$\times 4$
Urban100 ×$\times 4$
Urban100 ×$\times 6$
Manga109 ×$\times 6$
Manga109 ×$\times 6$
LR input LIIF [11] LTE [22] CiaoSR [8] Thera-L (ours) GT
Figure 10: Further qualitative examples for ×$\times 4$ and ×$\times 6$ upsampling factors. Best viewed zoomed in.
Appendix 0.F Computational Requirements
For comparison of computational requirements between methods, we provide training time, inference time, and inference VRAM requirements in Tab. 7, for the EDSR-baseline backbone. These values assume NVIDIA GeForce RTX 3090 Ti GPUs, and they show that we are more efficient than recent transformer-based methods, but still reach superior performance, as shown in the main paper.
Table 7: Computational requirements of various methods.
w/ EDSR-b Training time Inference time VRAM
(GPU days)
MetaSR 0.42 3.68 ms 588 MiB
LIIF 0.71 13.7 ms 1174 MiB
LTE 0.79 17.1 ms 1250 MiB
CiaoSR $\sim 2.6$ 37.7 ms 6474 MiB
CLIT $\sim 4.6$ 16.0 ms 2984 MiB
Thera-S 1.0 3.52 ms 754 MiB
Thera-M 2.2 10.0 ms 2290 MiB
Thera-L 3.3 14.3 ms 2292 MiB

### v3

@@ Appendix F Reproducibility of Existing Methods
We encountered challenges attempting to recreate the results reported for some of the competing methods, which explains why some are missing or differ from the originally reported numbers. Details are provided below.
@@
CUF [44]. At the time of writing, there were no public code or checkpoints available for CUF. Therefore, we could not generate numbers for datasets and scaling factors not reported in the original paper. We have denoted those missing values with “—” in the tables. Furthermore, we could not create any qualitative samples using CUF.
@@
CLIT [9]. For CLIT, at the time of writing, there is a public code, but no checkpoints. We have made a bona fide attempt to reproduce the models, but due to the cascaded training schedule and the large model size, the training process would require excessive amounts of compute: over a month using 8×8$\times$8 × Nvidia GeForce RTX 3090 GPUs. Unfortunately, the authors did not respond to our requests for the trained checkpoints used in their paper.
@@
CiaoSR [7]. We found that in the official CiaoSR implementation border cropping prior to evaluating on DIV2K deviated slightly from all other methods. We thus adapted the evaluation code to match the competition and enable a meaningful comparison. All DIV2K numbers were re-computed with the corrected code, resulting in slight deviations from those reported in the original paper.
MSIT [56]. The numbers reported in the paper were achieved by training on a roughly 3$\times$ larger training set, comprising not only DIV2K but also Flickr2K. For a fair comparison with all other ASR methods used in our paper, we re-trained MSIT on DIV2K alone. We used the authors’ official code and configuration files and trained for two stages, each comprising 1050 epochs. For the second (RiM) training stage we used scaling factors between ×$\times 1$ and ×$\times 4$. Performance generally drops as a result of the smaller training set, in line with the ablation experiments reported for MSIT, where the models were also trained only with DIV2K.